OpenStudy (anonymous):

a train travelling at a speed of 45 km/hr takes 10 seconds to pass a certain point . then the length of the train is

OpenStudy (anonymous):

45*5*10/18=125m

OpenStudy (dumbcow):

matricked converted speed to m/s 45*(1000/3600) = 45*(5/18) = 25/2 = 12.5 m/s multiply by 10 sec 12.5 *10 = 125 m